OpenStudy (anonymous):

Find the x- and y-intercepts and graph the equation plotting the intercepts. Please show all of your work.-4x - 5y=30

OpenStudy (anonymous):

-4x-5y=30 0-5y = 30 y = -6 (0,-6) -4x-0= 30 x = -30/4 (-7.5,0)